# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v19

@@ Abstract
As a rigorous extension of conventional Taylor expansion, statistical Taylor expansion replaces precise input variables with random input variables of known distributions and sample counts, to compute the mean, standard deviation, and reliability of each result. This approach tracks the propagation of input uncertainties through intermediate steps, making the final result path independent. Therefore, it differs fundamentally from conventional approaches that are usually path dependent, such that it may standardize numerical computations for analytic expressions. This study further presents an implementation of statistical Taylor expansion termed variance arithmetic, and demonstrates its performance across diverse mathematical applications.
@@ 1 Introduction
  • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [3][4][5], whereas statistical Taylor expansion extends this framework to include higher-order interactions as expressed in Formula (2.3) in this study.
@@ 2.2 Distributional Zero and Distributional Pole
  • A distributional pole occurs when $f^{(1)}_{x}=0\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=\infty$. For example, $(\mu\pm 1)^{2}$ has a distributional pole at $\tilde{x}=0$, as shown in Figure 1.
  • A distributional zero occurs when $f^{(1)}_{x}=\infty\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=0$. For example $\sqrt{\mu\pm 1}$ has a distributional zero at $\tilde{x}=0$, as shown in Figure 2.
@@ 2.3 Statistical Taylor Expansion
$\overline{f(x)}$ $=f(x)+\sum_{n=1}^{\infty}(\delta x)^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n,\kappa);$ (2.5)
$\delta^{2}f(x)$ $=\sum_{n=1}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.6)
@@
$\overline{xy}$ $=\zeta_{x}(0,\kappa_{x})x\;\zeta_{y}(0,\kappa_{y})y;$ (2.12)
$\delta^{2}(xy)$ $=\zeta_{x}(2,\kappa_{x})(\delta x)^{2}y^{2}+x^{2}\zeta_{y}(2,\kappa_{y})(\delta y)^{2}+\zeta_{x}(2,\kappa_{x})(\delta x)^{2}\;\zeta_{y}(2,\kappa_{y})(\delta y)^{2};$ (2.13)
@@ 2.4 One-Dimensional Examples
$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}(\delta x)^{n}\frac{\zeta(n,\kappa)}{n!};$ (2.14)
@@
[…]
$\overline{\sin(x)}=$ $\sum_{n=0}^{\infty}(\delta x)^{n}\eta(n,x)\frac{\zeta(n,\kappa)}{n!};$ (2.19)
$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\eta(j,x)\eta(n-j,x)\frac{\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)}{j!(n-j)!};$ (2.20)
@@ 2.6 Statistical Bounding
Figure 6: Ideal leakage $\varepsilon$ (y-axis) for varying sample count $\kappa_{s}$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=5$ for Gaussian input uncertainties. When $f(x)=x$, ideal leakage equals linear leakage $\omega$.
Figure 7: Ideal leakage $\varepsilon$ (y-axis) for varying sample count $\kappa_{s}$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) for Uniform input uncertainties. When $f(x)=x$, ideal leakage equals linear leakage $\omega$.
$f(x)-\overline{f(x)}$ $\simeq(\delta x)^{2}\zeta(2,\kappa)f_{x}^{(2)};$ (2.28)
$\delta^{2}f(x)$ $\simeq(\delta x)^{2}\zeta(2,\kappa)(f_{x}^{(1)})^{2};$ (2.29)
Figures (2.28) and (2.29) show that the first-order approximation of Formulas (2.5) and (2.6) both contain the term $(\delta x)^{2}\zeta(2,\kappa)$. As $\kappa\rightarrow+\infty$, $\zeta(2,\kappa)\rightarrow 1$ such that $\delta^{2}f(x)$ reaches a stable value but the convergence range of $\delta^{2}f(x)$ reduces toward zero for Formulas (2.17) and (2.22). As a trade-off, the choice of $\kappa$ to calculate the stable $f(x)-\overline{f(x)}$ and $\delta^{2}f(x)$ is defined as the ideal bound range $\hat{\kappa}$, and the result is the ideal uncertainty bias $\widehat{\overline{f(x)}}-f(x)$ and ideal variance $\hat{\delta^{2}}f(x)$. For a Uniform distribution, by definition $\hat{\kappa}=\sqrt{3}$. For a Gaussian distribution, according to the 5-$\sigma$ rule for determining the statistical significance of experimental result [1][2][4], $\hat{\kappa}=5$ by default.
Define linear leakage $\omega(\kappa)\equiv 1-\zeta(2,\kappa)$. When sampling from a distribution, the sample mean $\overline{x}$ and sample deviation $\delta x$ approach the distribution mean $\mu$ and distribution deviation $\sigma$ respectively as the sample count $N$ increases [4]. This yields the sample linear leakage $\omega(\kappa,N)$ for the interval $\overline{x}-\varrho\delta x,\overline{x}+\kappa\delta x]$, in contrast to the distributional linear leakage $\omega(\kappa)$ for the interval $\mu-\varrho\sigma,\mu+\kappa\sigma]$. Because $\omega(\kappa)\neq\omega(\kappa,N)$ for finite $N$, let $\omega(\kappa)=\omega(\kappa_{s},N)$, where $\kappa_{s}$ is the measuring bound range, and $\kappa(\kappa_{s},N)$ is the measured bound range, with the latter used in calculating $\zeta(n,\kappa)$.
@@
Define bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$. When the underlying distribution is Normal, Formula (2.30) presents the difference between the linear leakage $\omega(\kappa)$ and the bounding leakage $\epsilon(\kappa)$, whereas Formulas (2.31) and (2.32) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4]. Figure 4 shows that $\omega(\kappa_{s})<\omega(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\omega(\kappa_{s},N)=\omega(\kappa_{s})$ along the y-axis direction, and that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction. Figure 5 sections Figure 4 along the y-axis direction for $\kappa_{s}=4,5,6$, and shows that to reach the stable variance for a $\kappa_{s}$, the required sample count is $N\gtrsim 10^{3}$, whereas the stable resulting linear leakage $\omega$ is smaller for larger measuring bounding range $\kappa_{s}$. It also shows that linear leakage converges to 0 with increasing sample count $N$ faster when $\kappa_{s}$ is larger.
@@
Define ideal leakage as $\varepsilon\equiv 1-\delta^{2}f/\widehat{\delta^{2}}f\simeq\omega$, which quantifies the reliability of $\hat{\delta^{2}}f$ from the underlying distributions and the corresponding sample counts of all inputs, with Formula (2.33) as an example. Figure 6 compares ideal leakage $\varepsilon$ and linear leakage $\omega$ for the selected functions when the input uncertainty distribution is Gaussian, and shows that the example stable ideal leakages are less than $10^{-4}$ when $\hat{\kappa}=5$, which is small enough for most applications.
The unbiased bounding range $\kappa_{min},\kappa_{max}]$ of Uniform distribution is estimated as $1+\frac{1}{N}$ times of the minimal and maximal values in sampled data set. The bonding leakage $\epsilon$ is the percentage of $\kappa_{min},\kappa_{max}]$ outside the true bounding range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$. Because $\zeta(2n)$ for Uniform distribution increases much slower than $\zeta(2n,5)$ for Gaussian distribution, Figure 7 shows that for the selected functions, ideal leakage equals linear leakage for Uniform input uncertainty, and both decrease linearly with sample count $N$. The difference between Figures 6 and 7 shows the effect of input uncertainty distribution.
@@ 2.7 Dependency Tracing
$\delta^{2}(f+g)=$ $\;\delta^{2}f+\delta^{2}g+2(\overline{fg}-\overline{f}\overline{g});$ (2.34)
$\delta^{2}(c_{1}f+c_{0})=$ $\;c_{1}^{2}\delta^{2}f;$ (2.35)
When all inputs satisfy the uncorrelated uncertainty condition, statistical Taylor expansion traces dependencies through the intermediate steps. For example:
  • Formula (2.34) expresses $\delta^{2}(f+g)$, The dependency tracing of is illustrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.11).
@@
  • Formula (2.35) gives the variance of the linear transformation of a function, which can be applied to other formulas for more general dependency tracing.
@@ 3.1 Monotonic
Similar trends holds when the input uncertainty follows other distribution. Figure 9 shows the upper bound $P(x)$ for $(x\pm\delta x)^{c}$ when the input uncertainty distribution is Uniform. It has almost the same trend of upper bounds $P(x)$ increasing with $c$ as shown in Figure 9 but just with different y-scaling: $1/\sqrt{3}$ instead of by $1/5$.
@@ 3.2 Positive
Figure 10: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for Gaussian uncertainty. The corresponding resulting uncertainty bias and deviation are also shown (right y-axis). The x-axis is expressed in units of $\pi$.
Figure 11: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for uniform uncertainty. The corresponding resulting uncertainty bias and deviation are also shown (right y-axis). The x-axis is expressed in units of $\pi$.
In addition to convergence, the variance expansion may yield negative results when the expansion is terminated arbitrarily, as in Formula (2.20) for $sin(x\pm\delta x)$. Figure 10 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$ for Gaussian input uncertainty. Beyond this upper bound, the result deviation is no longer positive at some expansion order. Conceptually, the upper-bound for $\delta x$ should not be much more than $\pi/2$ to make the result meaningful, which is satisfied as shown in Figure 10. Similar trend holds when the input uncertainty is uniform but with larger upper bound $\delta x$, as shown in Figure 11.
@@ 3.4 Comparison
When the value difference is zero, the two imprecise values are considered equal. In statistics, such two values have a $50\%$ possibility of being either less than or greater to each other but zero probability of being exactly equal [4]. In variance arithmetic, however, they are treated as neither less nor greater than each other and therefore are considered equal.
Otherwise, the standard z-statistic method [4] is applied to determine whether two imprecise values are statistically equal, less than, or greater than each other. For example, the difference between $1.002\pm 0.001$ and $1.000\pm 0.002$ is $0.002\pm 0.00224$, yielding $z=0.002/0.00224$. The probability that they are not equal is $\xi(|z|/\sqrt{2})=62.8\%$, in which $\xi(z)$ is the cumulative distribution function for Normal distribution [4]. If the threshold probability for inequality is set at $50\%$, then $1.000\pm 0.002<1.002\pm 0.001$. Alternatively, an equivalent bounding range for z can be used, such as $|z|\leq 0.67448975$ for an equal probability threshold of $50\%$.
Because the result of comparison depends on threshold probability which is application specific, comparison is not part of variance arithmetic.
@@ 4 Verification of Variance Arithmetic
When an input contains unspecified errors, such as numerical errors in library functions or floating-point rounding errors, Gaussian noise with progressively increasing deviations can be added to inputs, until ideal coverage is attained. The minimal noise deviation required provides a good estimate of the magnitude of the unspecified input uncertainty deviations. Achieving ideal coverage serves as a necessary verification step to ensure that statistical Taylor expansion has been applied correctly within the given context. The input noise range that yields ideal coverage defines the ideal application range for the analytic function.

### v20

@@ Abstract
As a rigorous extension of conventional Taylor expansion, statistical Taylor expansion replaces precise input variables with random input variables of known distributions and sample counts, to compute the mean, deviation, and reliability of each result. This approach tracks the propagation of input uncertainties through intermediate steps, making the final result path independent. Therefore, it differs fundamentally from conventional approaches that are usually path dependent, such that it may standardize numerical computations for analytic expressions. This study further presents an implementation of statistical Taylor expansion termed variance arithmetic, and demonstrates its performance across diverse mathematical applications.
@@ 1 Introduction
  • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [3][4][5], whereas statistical Taylor expansion extends this framework to include higher-order interactions as expressed in Formula (2.3) in this study. A full expansion illuminates the convergence of the result.
@@ 2.2 Distributional Zero and Distributional Pole
  • A distributional pole occurs when $y^{(1)}_{x}=0\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=\infty$. For example, $(\mu\pm 1)^{2}$ has a distributional pole at $\tilde{x}=0$, as shown in Figure 1.
  • A distributional zero occurs when $y^{(1)}_{x}=\infty\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=0$. For example $\sqrt{\mu\pm 1}$ has a distributional zero at $\tilde{x}=0$, as shown in Figure 2.
@@ 2.3 Statistical Taylor Expansion
$\overline{f(x)}$ $=f(x)+\sum_{n=2}^{\infty}(\delta x)^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n,\kappa);$ (2.5)
$\delta^{2}f(x)$ $=\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.6)
@@
$\frac{\overline{xy}}{xy}$ $=\zeta_{x}(0,\kappa_{x})\zeta_{y}(0,\kappa_{y});$ (2.12)
$P(xy)^{2}$ $=\zeta_{x}(2,\kappa_{x})P(x)^{2}+\zeta_{y}(2,\kappa_{y})P(y)^{2}+\zeta_{x}(2,\kappa_{x})P(x)^{2}\zeta_{y}(2,\kappa_{y})P(y)^{2};$ (2.13)
@@ 2.4 One-Dimensional Examples
$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=2}^{\infty}(\delta x)^{n}\frac{\zeta(n,\kappa)}{n!};$ (2.14)
@@
$\overline{\sin(x)}=$ $\sum_{n=0}^{\infty}(\delta x)^{n}\sin(x)_{x}^{(n)}\frac{\zeta(n,\kappa)}{n!};$ (2.18)
$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\sin(x)_{x}^{(j)}\sin(x)_{x}^{(n-j)}\frac{\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)}{j!(n-j)!};$ (2.19)
@@ 2.6 Statistical Bounding
Figure 6: Ideal leakage $\varepsilon$ (y-axis) for varying sample count NN (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=5$ for Gaussian input uncertainties. When $f(x)=x$, ideal leakage equals linear leakage $\omega$.
Figure 7: Ideal leakage $\varepsilon$ (y-axis) for varying sample count NN (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) for Uniform input uncertainties. When $f(x)=x$, ideal leakage equals linear leakage $\omega$.
The first-order approximation of Formulas (2.5) and (2.6) both contain the term $(\delta x)^{2}\zeta(2,\kappa)$. As $\kappa\rightarrow+\infty$, $\zeta(2,\kappa)\rightarrow 1$ such that $\delta^{2}f(x)$ reaches a stable value but the convergence range of $\delta^{2}f(x)$ reduces toward zero for Formulas (2.17) and (2.21). As a trade-off, the choice of $\kappa$ to calculate the stable $f(x)-\overline{f(x)}$ and $\delta^{2}f(x)$ is defined as the ideal bound range $\hat{\kappa}$, and the result is the ideal variance $\hat{\delta^{2}}f(x)$. For a Uniform distribution, by definition $\hat{\kappa}=\sqrt{3}$. For a Gaussian distribution, according to the 5-$\sigma$ rule for determining the statistical significance of experimental result [1][2][4], $\hat{\kappa}=5$ by default.
Define linear leakage $\omega(\kappa)\equiv 1-\zeta(2,\kappa)$. In $\delta^{x}x=\zeta(2,\kappa)(\delta x)^{2}$, the result variance $\delta^{x}$ is less than input variance $(\delta x)^{2}$ due to linear leakage. When sampling from a distribution, the sample mean $\overline{x}$ and sample deviation $\delta x$ approach the distribution mean $\mu$ and distribution deviation $\sigma$ respectively as the sample count $N$ increases [4]. This yields the sample linear leakage $\omega(\kappa,N)$ for the interval $\overline{x}-\varrho\delta x,\overline{x}+\kappa\delta x]$, in contrast to the distributional linear leakage $\omega(\kappa)$ for the interval $\mu-\varrho\sigma,\mu+\kappa\sigma]$. Because $\omega(\kappa)\neq\omega(\kappa,N)$ for finite $N$, let $\omega(\kappa)=\omega(\kappa_{s},N)$, where $\kappa_{s}$ is the measuring bound range, and $\kappa(\kappa_{s},N)$ is the measured bound range, with the latter used in calculating $\zeta(n,\kappa)$.
@@
Define bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$. When the underlying distribution is Normal, Formula (2.27) presents the difference between the linear leakage $\omega(\kappa)$ and the bounding leakage $\epsilon(\kappa)$, whereas Formulas (2.28) and (2.29) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4]. Figure 4 shows that $\omega(\kappa_{s},N+1)<\omega(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\omega(\kappa_{s},N)=\omega(\kappa_{s})$ along the y-axis direction, and that $\kappa(\kappa_{s},N+1)<\kappa(\kappa_{s},N)$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction. Figure 5 sections Figure 4 along the y-axis direction for $\kappa_{s}=4,5,6$, and shows that to reach the stable variance for a $\kappa_{s}$, the required sample count is $N\gtrsim 10^{3}$, whereas the stable resulting linear leakage $\omega$ is smaller for larger measuring bounding range $\kappa_{s}$. It also shows that linear leakage converges toward 0 with increasing sample count $N$ faster when $\kappa_{s}$ is larger.
@@
$\varepsilon(xy)$ $=\frac{2\omega_{x}P(x)^{2}+2\omega_{y}P(y)^{2}-\omega_{x}P(x)^{2}\omega_{y}P(y)^{2}}{P(x)^{2}+P(y)^{2}+P(x)^{2}P(y)^{2}};$ (2.31)
Define ideal leakage as $\varepsilon\equiv 1-\delta^{2}f/\widehat{\delta^{2}}f\simeq\omega$, which quantifies the reliability of $\hat{\delta^{2}}f$ from the underlying distributions and the corresponding sample counts of all inputs, with Formulas (2.30) and 2.31 as two examples. Figure 6 compares ideal leakage $\varepsilon$ and linear leakage $\omega$ for the selected functions when the input uncertainty distribution is Gaussian, and shows that the example stable ideal leakages are less than $10^{-4}$ when $\hat{\kappa}=5$, which is small enough for most applications.
The unbiased bounding range $\kappa_{min},\kappa_{max}]$ of Uniform distribution is estimated as $1+\frac{1}{N}$ times of the minimal and maximal values in sampled data set [20]. The bounding leakage $\epsilon$ is the percentage of $\kappa_{min},\kappa_{max}]$ outside the true bounding range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$. Because $\zeta(2n)$ for Uniform distribution increases much slower than $\zeta(2n,5)$ for Gaussian distribution, Figure 7 shows that for the selected functions, ideal leakage equals linear leakage for Uniform input uncertainty, and both decrease linearly with sample count $N$. The difference between Figures 6 and 7 shows the effect of input uncertainty distribution to the output deviation $\delta f$.
@@ 2.7 Dependency Tracing
When all inputs satisfy the uncorrelated uncertainty condition, statistical Taylor expansion traces dependencies through the intermediate steps. For example, when the sample count $N$ is sufficiently large:
  • The dependency tracing of $\delta^{2}(f+g)$ is illustrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.11).
@@ 3.1 Monotonic
Similar trends holds when the input uncertainty follows other distribution. Figure 9 shows the upper bound $P(x)$ for $(x\pm\delta x)^{c}$ when the input uncertainty distribution is Uniform. It has almost the same trend of upper bounds $P(x)$ increasing with $c$ as shown in Figure 9 but just with different y-scaling: $1/\sqrt{3}$ instead of $1/5$.
@@ 3.2 Positive
Figure 10: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for Gaussian uncertainty. The corresponding resulting uncertainty bias and deviation are also shown (right y-axis).
Figure 11: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for uniform uncertainty. The corresponding resulting uncertainty bias and deviation are also shown (right y-axis).
In addition to convergence, the variance expansion may yield negative results when the expansion is terminated arbitrarily, as in Formula (2.19) for $sin(x\pm\delta x)$, such that an upper bound for $\delta x$ is need to guarantee non-negative variance. Figure 10 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$ for Gaussian input uncertainty. Because $\sin(x)\in[-1,+1]$, $\overline{\sin(x)}\pm\delta\sin(x)$ should also follow this range, as indicated in 10. Similar trend holds when the input uncertainty is uniform but with larger upper bound $\delta x$, as shown in Figure 11.
@@ 3.4 Comparison
When the value difference is zero, the two imprecise values are considered equal. In statistics, such two values have a $50\%$ possibility of being either less than or greater to each other but zero probability of being exactly equal [4]. In variance arithmetic, however, they are considered equal.
Otherwise, the standard z-statistic method [4] is applied to determine whether two imprecise values are statistically equal, less than, or greater than each other. For example, the difference between $1.002\pm 0.001$ and $1.000\pm 0.002$ is $0.002\pm 0.00224$, yielding $z=0.002/0.00224$. The probability that they are not equal is $\xi(|z|/\sqrt{2})=62.8\%$, in which $\xi(z)$ is the cumulative distribution function for Normal distribution [4]. If the threshold probability for inequality is set at $50\%$, then $1.000\pm 0.002<1.002\pm 0.001$. Because the result of comparison depends on threshold probability which is application specific, comparison is not part of variance arithmetic.
@@ 4 Verification of Variance Arithmetic
When an input contains unspecified errors, such as numerical errors in library functions or floating-point rounding errors, Gaussian noise with progressively increasing deviations can be added to inputs, until ideal coverage is attained. The minimal noise deviation required for the ideal coverage provides a good estimate of the magnitude of the unspecified input uncertainty deviations. Achieving ideal coverage serves as a necessary verification step to ensure that statistical Taylor expansion has been applied correctly within the given context. The input noise range that yields ideal coverage defines the ideal application range for the analytic function.